# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

### v1

@@ Abstract
Sparse Autoencoders (SAEs) extract features from LLM internal activations, meant to correspond to single concepts. A core SAE training hyperparameter is L0: how many features should fire per token on average. Existing work compares SAE algorithms using sparsity–reconstruction tradeoff plots, implying L0 is a free parameter with no single correct value. In this work we study the effect of L0 on BatchTopK SAEs, and show that if L0 is not set precisely, the SAE fails to learn the underlying features of the LLM. If L0 is too low, the SAE will mix correlated features to improve reconstruction. If L0 is too high, the SAE finds degenerate solutions that also mix features. Further, we demonstrate a method to determine the correct L0 value for an SAE on a given training distribution, which finds the true L0 in toy models and coincides with peak sparse probing performance in LLMs. We find that most commonly used SAEs have an L0 that is too low. Our work shows that, to train SAEs with correct features, practitioners must set L0 correctly.
@@ 1 Introduction
Sparse autoencoders (SAEs) decompose the dense, polysemantic activations of LLMs into interpretable latent features (Cunningham et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen and Field, 1997). SAEs have the advantage of being unsupervised, and can be scaled to millions of neurons in its hidden layer (hereafter called “latents”) (Templeton et al., 2024; Gao et al., 2024). When training an SAE, practitioners must decide on the L0 of the SAE: that is, the sparsity; how many latents activate on average for a given input. [^1] TopK and BatchTopK SAEs (Gao et al., 2024; Bussmann et al., 2024) set the L0 ($K$) directly, whereas L1 and JumpReLU (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024) adjust it via a coefficient in the loss. In any case, all SAE trainers must decide on the target L0. L0 is typically considered a neutral design choice: most of the literature evaluates SAEs at a range of L0 values, referring to this as a “sparsity–reconstruction tradeoff” (Gao et al., 2024; Rajamanoharan et al., 2024), implying any L0 is equally valid.
@@
However, recent work shows the same trend: too low an L0 leads to worse SAE performance on downstream tasks Kantamneni et al. (2025); Bussmann et al. (2025). In this work, we explore the effect of L0 on SAEs. Starting by exploring a toy model with correlated features, we demonstrate that if the L0 is too low, the SAE can “cheat” by mixing together components of correlated features, achieving better reconstruction compared to an SAE with correct, disentangled features. We consider this to be a manifestation of feature hedging (Chanin et al., 2025), where the SAE abuses feature correlations to compensate for insufficient resources to correctly model the underlying features. Furthermore, we show that if the L0 is too high, the SAE can learn degenerate mixtures of features. We show that it is possible to determine the correct L0 of an SAE by observing projection magnitude between the SAE decoder and training activations.
@@
We validate these findings on Gemma-2-2b (Team et al., 2024), demonstrating that the same decoder patterns we observe in our toy model experiments also manifest in LLM SAEs. We further validate that the optimal L0 we find with our method in Gemma-2-2b matches with peak performance on sparse probing tasks (Kantamneni et al., 2025). Our findings show that L0 must be set correctly for SAEs to learn correct features, and implies that most SAEs used by researchers today have too low an L0.
@@ 2 Background
An SAE decomposes an input activation $\mathbf{x}\in\mathbb{R}^{d}$ into a hidden state $\mathbf{a}$ consisting of $h$ hidden neurons, called “latents”. An SAE is composed of an encoder $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{h\times d}$, a decoder $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times h}$, a decoder bias $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$, and encoder bias $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{h}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2024) or BatchTopK (Bussmann et al., 2024).
$\mathbf{a}=$ $\sigma(\mathbf{W}_{\text{enc}}(\mathbf{x}-\mathbf{b}_{\text{dec}})+\mathbf{b}_{\text{enc}})$ (1)
$\hat{\mathbf{x}}$= $\mathbf{W}_{\text{dec}}\mathbf{a}+\mathbf{b}_{\text{dec}}$ (2)
@@ 3 Toy model experiments
We set up a toy model with 50 mutually orthogonal true features $F=\{f_{0},\dots,f_{49}\}\in\mathbb{R}^{100}$. Each feature $f_{i}$ has firing probability $P_{i}$. We set $P(f_{0})=0.345$, and we linearly decrease $P_{i}$ to $P_{49}=0.05$, so feature firing probability decreases with feature number. We randomly generate a correlation matrix, so the firings of each feature are correlated with other features. Features fire with magnitude $\sim\mathcal{N}(1.0,0.15)$. We sampling true feature firings, and summing all firing features to create a training input for the SAE. The goal of the SAE, then, is to learn these true features despite only being trained on their summed activations. Since we have ground-truth knowledge, we know the true number of features firing on average. We call this the true L0, which is 11 for this toy model.
Throughout this work, we use BatchTopK SAEs (Bussmann et al., 2024), as this architecture allows directly controlling L0 rather than controlling it indirectly via a L1 coefficient (Cunningham et al., 2024; Bricken et al., 2023) or L0 coefficient (Rajamanoharan et al., 2024) in the loss. For these toy model experiments, we train SAEs on 15M synthetic samples using SAELens (Bloom et al., 2024).
(a) SAE L0 = 5
(b) SAE L0 = 11
(c) SAE L0 = 18
Figure 1: SAE decoder cosine similarity with true features. True features L0 = 11.
@@
We train SAEs with L0 that is too low (L0=5), exactly correct (L0=11), and too high (L0=18). Results are shown in Figure 1. When the SAE L0 matches the true L0 (Figure 1(b)), the SAE exactly learns the true features. When the SAE L0 is too low (Figure 1(a)), the SAE mixes components of correlated features together, especially breaking latents tracking high-frequency features. When SAE L0 is too high (Figure 1(c)), the SAE learns degenerate solutions that mix features together. The further the L0 is from the true L0, the more broken the SAE becomes. Interestingly, when L0 is too high the SAE still learns many correct latents, but when L0 is too low, every latent in the SAE is affected.
3.1 MSE loss incentivizes low-L0 SAEs to learn incorrect features
@@
Why do SAEs with too low L0 not learn the true features? We take the correct SAE from Figure 1(b) and set L0=5, to match the SAE from Figure 1(a). We then generate 100k training samples and calculate the Mean Square Error (MSE) of both these SAEs. The SAE with incorrect latents from Figure 1(a) achieves a MSE of $2.73$, while the SAE with ground-truth correct latents achieves a much worse MSE of $4.88$. Thus, MSE loss actively incentivizes low L0 SAEs to learn incorrect latents. Furthermore, this shows that using reconstruction error as a target, and thus the sparsity–reconstruction tradeoff (Gao et al., 2024; Rajamanoharan et al., 2024; Cunningham et al., 2024) is harmful: when the L0 is too low, the incorrect SAE achieves better reconstruction than the ground-truth correct SAE. We explore this further in Appendix A.2.
@@ 3.2 Detecting the true L0 using the SAE decoder
Figure 1 reveals that the SAE decoder latents contain mixes of underlying features, both when the L0 is too high and also when it is too low. As the SAE approaches the correct L0, each SAE latent has fewer components of multiple true features mixed in, becoming more monosemantic. Thus, we expect that when the SAE has the correct L0, most latents should have near zero projection on arbitrary training inputs, because they usually do not contain the feature being tracked by that latent. If we are far from the correct L0, then SAE latents contain components of many underlying features, and we expect latents to project unexpectedly strongly on arbitrary training inputs.
We now define a metric we call Nth decoder projection score, or $s_{n}^{\text{dec}}$, that we can minimize to find the optimal L0 of the SAE. Given SAE inputs $\mathbf{x}\in\mathbb{R}^{b\times d}$ where $b$ is the batch size and $d$ is the input dimension, we first compute the decoder projections for all latents:
$\mathbf{Z}=(\mathbf{x}-\mathbf{b}_{\text{dec}})\mathbf{W}_{\text{dec}}^{\top}\in\mathbb{R}^{b\times h}$ (3)
where $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$ is the decoder bias and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times h}$ is the decoder weight matrix with $h$ latent dimensions. To aggregate across the batch, we flatten $\mathbf{Z}$ to obtain $\mathbf{z}\in\mathbb{R}^{bh}$ and sort these values in descending order to get $\mathbf{z}_{\downarrow}$. The Nth decoder projection is then defined as:
$s_{n}^{\text{dec}}=\mathbf{z}_{\downarrow}[n\cdot b]$ (4)
@@
Figure 2: Nth decoder projection vs SAE L0 for N=12 (left) and N=20 (right) on our toy model SAEs. The true L0, 11, is marked by a dotted line on the plots. Both settings of N are minimized at the true L0, although the slopes of the metric are slightly different depending on N.
We calculate $s_{n}^{\text{dec}}$ for $n=18$ and $n=22$, with SAE L0 ranging from 2 to 25 in Figure 2. We see that the metric is minimized at the true L0, 11, in both cases, although the slope of the metric changes depending on $n$. In both cases, the slope of $s_{n}^{\text{dec}}$ is flat when L0 is slightly higher than the true L0.

### v2

@@ Abstract
Sparse Autoencoders (SAEs) extract features from LLM internal activations, meant to correspond to interpretable concepts. A core SAE training hyperparameter is L0: how many SAE features should fire per token on average. Existing work compares SAE algorithms using sparsity–reconstruction tradeoff plots, implying L0 is a free parameter with no single correct value aside from its effect on reconstruction. In this work we study the effect of L0 on SAEs, and show that if L0 is not set correctly, the SAE fails to disentangle the underlying features of the LLM. If L0 is too low, the SAE will mix correlated features to improve reconstruction. If L0 is too high, the SAE finds degenerate solutions that also mix features. Further, we present a proxy metric that can help guide the search for the correct L0 for an SAE on a given training distribution. We show that our method finds the correct L0 in toy models and coincides with peak sparse probing performance in LLM SAEs. We find that most commonly used SAEs have an L0 that is too low. Our work shows that L0 must be set correctly to train SAEs with correct features.
@@ 1 Introduction
It is theorized that Large Language Models (LLMs) represent concepts as linear directions in representation space, known as the Linear Representation Hypothesis (LRH) (Elhage et al., 2022; Park et al., 2024). These concepts are nearly orthogonal linear directions, allowing the LLM to represent many more concepts than there are neurons, a phenomenon known as superposition (Elhage et al., 2022). However, superposition poses a challenge for interpretability, as neurons in the LLM are polysemantic, firing on many different concepts.
@@
Sparse autoencoders (SAEs) are meant to reverse superposition, and extract interpretable, monosemantic latent features (Cunningham et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen & Field, 1997). SAEs have the advantage of being unsupervised, and can be scaled to millions of neurons in its hidden layer (hereafter called “latents” [^1] We use latents to prevent overloading the term feature, which we reserve for human-interpretable concepts the SAE may capture. This breaks from earlier usage which used feature for both (Elhage et al., 2022), but aligns with the terminology in (Lieberum et al., 2024) and makes the distinction more clear.). When training an SAE, practitioners must decide on the sparsity of SAE, measured in terms of L0, or how many latents activate on average for a given input.  [^2] TopK and BatchTopK SAEs (Gao et al., 2024; Bussmann et al., 2024) set the L0 ($K$) directly, whereas L1 and JumpReLU (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024) adjust it via a coefficient in the loss. In any case, all SAE trainers must decide on the target L0. L0 is typically considered a neutral design choice: most of the literature evaluates SAEs at a range of L0 values, referring to this as a “sparsity–reconstruction tradeoff” (Gao et al., 2024; Rajamanoharan et al., 2024). While most practitioners would expect that too high an L0 will break the SAE (afterall, it is called a sparse autoencoder), the implication of “sparsity–reconstruction tradeoff” plots is that any sufficiently low L0 is equally valid.
@@
However, recent work shows the same trend: low L0 SAEs perform worse on downstream tasks Kantamneni et al. (2025); Bussmann et al. (2025). What causes this degraded performance at low L0? In this work, we explore the effect of L0 on SAEs. We begin with toy model experiments using synthetic data, and show that if the L0 is too low, the SAE can “cheat” by mixing together components of correlated features, achieving better reconstruction compared to an SAE with correct, disentangled features. We consider this to be a manifestation of feature hedging (Chanin et al., 2025), where the SAE abuses feature correlations to compensate for insufficient resources to model the underlying features monsemantically. This mixing of correlated features into SAE latents affects both positively and negatively correlated features, meaning that in low L0 SAEs, nearly all latents are both less interpretable and more noisy than an SAE with a correctly set L0.
Our findings also show that “sparsity–reconstruction tradeoff” plots, commonly used to assess SAE architectures, are not a sound method of evaluating SAEs. We demonstrate using toy model experiments that at low L0, an SAE with ground-truth correct latents achieves worse reconstruction than an SAE that mixes correlated features. Thus, if we had an SAE training method that resulted in perfect SAEs, “sparsity–reconstruction tradeoff” plots would cause us to reject that method.
@@
Finally, we develop a proxy metric based on projections between the SAE decoder and training activations that can detect if L0 is too low. We validate these findings on Gemma-2-2b (Team et al., 2024), demonstrating that decoder patterns similar to what we observe in our toy model experiments also manifests in LLM SAEs. We further validate that the optimal L0 we find with our method in Gemma-2-2b matches peak performance on sparse probing tasks (Kantamneni et al., 2025).
Our findings are of direct importance to anyone using SAEs in practice, showing that L0 must be set correctly for SAEs to learn correct features. Furthermore, our work implies that most SAEs used by researchers today have too low an L0.
Figure 1: When SAE L0 is too low (left) or too high (right), the SAE mixes together correlated features, ruining monsemanticity. Only at the correct L0 (middle), the SAE learns correct features.
@@ 2 Background
An SAE decomposes an input activation $\mathbf{x}\in\mathbb{R}^{d}$ into a hidden state, $\mathbf{a}$, consisting of $h$ hidden neurons, called “latents”. An SAE is composed of an encoder $\mathbf{W}_{\text{\text{enc}}}\in\mathbb{R}^{h\times d}$, a decoder $\mathbf{W}_{\text{\text{dec}}}\in\mathbb{R}^{d\times h}$, a decoder bias $\mathbf{b}_{\text{\text{dec}}}\in\mathbb{R}^{d}$, and encoder bias $\mathbf{b}_{\text{\text{enc}}}\in\mathbb{R}^{h}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2024) or BatchTopK (Bussmann et al., 2024). The decoder is sometimes called the dictionary, in reference to sparse dictionary learning. We use both terms interchangeably.
$\mathbf{a}=$ $\sigma(\mathbf{W}_{\text{\text{enc}}}(\mathbf{x}-\mathbf{b}_{\text{\text{dec}}})+\mathbf{b}_{\text{\text{enc}}})$ (1)
$\hat{\mathbf{x}}$= $\mathbf{W}_{\text{\text{dec}}}\mathbf{a}+\mathbf{b}_{\text{\text{dec}}}$ (2)
In this work we focus on BatchTopK and JumpReLU SAEs as these are both considered SOTA architectures. The JumpReLU activation is a modified ReLU with a threshold parameter $\tau>0$, so $\text{JumpReLU}_{\tau}(x)=x\cdot\mathbf{1}_{x>\tau}$. The BatchTopK activation function selects the top $b\times k$ activations across a batch of size $b$, allowing variance in the $k$ selected per sample in the batch. After training, a BatchTopK SAE is converted to a JumpReLU SAE with a global $\tau$. We follow the JumpReLU training procedure outlined by Anthropic (Conerly et al., 2025).
SAEs are trained as follows, with an auxiliary loss $\mathcal{L}_{p}$ to revive dead latents with corresponding coefficient $\lambda_{p}$. JumpReLU SAEs also have a sparsity loss $\mathcal{L}_{s}$ and corresponding coefficient $\lambda_{s}$.
$\mathcal{L}=\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert_{2}^{2}+\lambda_{s}\mathcal{L}_{s}+\lambda_{p}\mathcal{L}_{p}$ (3)
The formulation of $\mathcal{L}_{s}$ and $\mathcal{L}_{p}$ for JumpReLU and BatchTopK SAEs is shown in Appendix A.1.
@@ 3 Toy model experiments
The Linear Representation Hypothesis (LRH) (Elhage et al., 2022; Park et al., 2024) states that LLMs represent concepts (alternatively referred to as “features”) as (nearly) orthogonal linear directions in representation space. Thus, the hidden activations in an LLM are simply the sum all the firing feature vectors (a feature direction with a positive, non-zero magnitude) that are being represented. While an LLM can represent a potentially large number of concepts this way, in any given activation, only a small number of concepts are actively represented.
For instance, if we inspect a hidden activation from within an LLM at the token “_Canada”, we may expect this activation to be a sum of feature vectors representing concepts like “country”, “North America”, “starts with C”, “noun”, etc… The job of a sparse autoencoder is to recover these “true feature” directions in its dictionary.
In a real LLM, we do not have ground-truth knowledge of the “true features” the model is representing, so we do not know if the SAE has learned the correct features. Fortunately, it is easy to create a toy model setup that follows the requirements of the LRH while providing ground-truth knowledge of the underlying true features.
Our toy model has a set of feature embeddings $\mathbf{F}\in\mathbb{R}^{g\times d}$, where $d$ is the input dimension of our SAE, and $g$ is the number of features. All features are orthogonal, so $f_{i}\cdot f_{j}=0$ for $i\neq j$. Each feature $f_{i}$ fires with probability $p_{i}$, mean magnitude $\mu_{i}$, and magnitude standard deviation $\sigma_{i}$. Feature activations follow a correlated Bernoulli process controlled by correlation matrix $C$, with final magnitudes given by $m_{i}=a_{i}\cdot\text{ReLU}(\mu_{i}+\sigma_{i}\epsilon_{i})$, where $a_{i}$ indicates whether feature $i$ is active and $\epsilon_{i}\sim\mathcal{N}(0,1)$. Training activations for an SAE, $x\in\mathbb{R}^{d}$, are thus generated as $x=\sum_{i=1}^{n}m_{i}f_{i}$
In these toy model experiments, we mainly focus on BatchTopK SAEs (Bussmann et al., 2024) as this enables direct control of L0. Additionally, we validate our results with JumpReLU SAEs. We train SAEs on 15M synthetic samples with batch size 500 using SAELens (Bloom et al., 2024).
Throughout this section we will use the following terminology:
True L0
In toy models we have complete control over which features fire, so we know how many features are firing on average. We refer to this as the true L0 of the toy model.
Ground-truth SAE
Since we know the ground-truth features in our toy models, we can construct an SAE that perfectly captures these features. We refer to this as the ground-truth SAE. This is an SAE where $g=h$, $\mathbf{W}_{\text{enc}}=\mathbf{F}^{T}$, $\mathbf{W}_{\text{dec}}=\mathbf{F}$, $\mathbf{b}_{\text{enc}}=0$, $\mathbf{b}_{\text{dec}}=0$.
3.1 Low L0 SAEs mix correlated and anti-correlated features
We begin with a small toy model with 5 true features ($g=5$) in an input space of $d=20$. We set each $p_{i}=0.4$ such that on average $2$ features are active per input, for a true L0 of 2. We begin with a simple correlation pattern between features, where $f_{0}$ is positively correlated with every feature $f_{1}$ through $f_{4}$, but otherwise there are no other correlations. We then train an SAE with $L0=2$, matching the true L0 of the model, and an SAE with slightly lower value of $L0=1.8$ (BatchTopK SAEs permit setting fractional L0). For the $L0=1.8$ SAE, we initialize it to the ground-truth solution, ensuring that the result of training is due to gradient pressure rather than just being a local minimum. We show the toy model feature correlation matrix as well as decoder cosine similarity plots with the true features for both SAEs in Figure 2.
When the SAE L0 matches the true L0, we see that the SAE perfectly learns the underlying true features. However, when SAE L0 is smaller than the true L0, the resulting SAE latents mix feature components together based on the correlation matrix. The latents tracking features $f_{1}$ through $f_{4}$ all mix in a positive component of $f_{0}$, but they have no components of each other.
Next, we invert the correlation, i.e. each feature $f_{1}$ through $f_{4}$ is negatively correlated with $f_{0}$ instead, while keeping everything else unchanged. We show the correlation matrix and SAE decoder cosine similarity with true features plots in Figure 3.
Now, we see the same pattern as with positive correlations except inverted. The latents tracking features $f_{1}$ through $f_{4}$ mix in a negative component of $f_{0}$, but have no component of each other.
This pattern is problematic because it means that if our L0 is too low, every SAE latent will contain positive components of every positively correlated feature, and negative components of every negatively correlated feature in the model. Negative correlations are particularly bad, as negative correlations are prevalent throughout language. For instance, we may expect a nonsensical negative component of “Harry Potter” to appear in the latent for “French poetry”, since Harry Potter has nothing to do with French poetry. This will result in highly polysemantic and noisy SAE latents.
Figure 2: (left) Toy model feature correlation matrix showing positive correlations between features. (middle) SAE decoder cosine similarities with true feature when SAE L0 = 2, matching the true L0 of the toy model. (right) SAE decoder cosine similarities with true features when SAE L0 = 1.8. When L0 is too low, the SAE mixes components of features based on their firing correlations.
Figure 3: (left) Toy model feature correlation matrix showing negative correlations between features. (middle) SAE decoder cosine similarities with true feature when SAE L0 = 2, matching the true L0 of the toy model. (right) SAE decoder cosine similarities with true features when SAE L0 = 1.8,. When L0 is too low, the SAE mixes negative components of anti-correlated features.
3.2 Larger toy model experiments
Next, we scale up to a larger toy model with 50 true features ($g=50$) in input space of $d=100$. We set $p_{0}=0.345$ and linearly decrease to $p_{49}=0.05$, so firing probability decreases with feature number. The true L0 of this model is 11. We randomly generate a correlation matrix, so the firings of each feature are correlated with other features. Feature correlations are shown in Appendix A.2.
@@ 3.2 Larger toy model experiments
We train SAEs with L0 values that are too small ($L0=5$), exactly correct ($L0=11$), and too large ($L0=18$). Results are shown in Figure 1. When the SAE L0 matches the true L0, the SAE exactly learns the true features. When SAE L0 is too low, the SAE mixes components of correlated features together, particularly breaking latents tracking high-frequency features. When L0 is too high, the SAE learns degenerate solutions that mix features together. The further SAE L0 is from the true L0, the worse the SAE. Interestingly, when L0 is too high the SAE still learns many correct latents, but when L0 is too low, every latent in the SAE is affected.
3.3 MSE loss incentivizes low-L0 SAEs to mix correlated features
@@
Why do SAEs with low L0 not learn the true features? We construct a ground-truth SAE and set $L0=5$, to match the low L0 SAE from Figure 1. We generate 100k synthetic training samples and calculate the Mean Square Error (MSE) of both these SAEs. The trained SAE with incorrect latents achieves a MSE of $2.73$, while the ground-truth SAE achieves a much worse MSE of $4.88$. Thus, MSE loss actively incentivizes low L0 SAEs to learn incorrect latents.
3.4 The sparsity–reconstruction tradeoff
It is common practice to evaluate SAE architectures using a sparsity–reconstruction tradeoff plot (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), where the assumption is that having better reconstruction at a given sparsity is inherently better, and indicates that the SAE is correct. Afterall, we train SAEs to reconstruct inputs, so surely an SAE that has better reconstruction must therefore be a better SAE than one that has lower reconstruction?
Sadly, this is not the case. As we discussed in Section 3.3, when the L0 of the SAE is lower than optimal, the SAE can find ways to “cheat” by engaging in feature hedging (Chanin et al., 2025), and get a better MSE score by mixing components of correlated features together. This results in an SAE where the latents are not monosemantic, and do not track ground-truth features.
Figure 4: Sparsity ($L0$, lower is better) vs reconstruction (variance explained, higher is better) for learned SAEs and a ground-truth SAE. When L0 is less than the true L0 of the toy model (the dotted line), the trained SAE gets better reconstruction than the ground-truth SAE. Sparsity–reconstruction plots like this lead us to the incorrect conclusion that the ground-truth SAE is a worse SAE.
We next explore the sparsity–reconstruction tradeoff by training SAEs on our toy model at various L0s. Since we know the ground-truth features in our toy model, we construct a ground-truth SAE that perfectly represents these features. We vary the L0 of the ground truth SAE while leaving the encoder and decoder fixed at the correct features. We plot the variance explained vs L0 in Figure 4 for both SAEs. When the SAE L0 is lower than the true L0 of the toy model, the ground-truth SAE scores worse on reconstruction than the trained SAE! If we had an SAE training technique that gave us the ground truth correct SAE for a given LLM, sparsity–reconstruction plots would cause us to discard the correct SAE in favor of an incorrect SAE that mixes features together.
Figure 5: SAE decoder cosine similarity with true features for the learned SAEs with L0=1 (left) and L0=2 (middle), compared with the ground-truth SAE (right). The learned SAEs score much better than the ground truth SAE on variance explained, despite their corrupted, polysemantic latents.
We show the cosine similarity of the SAE decoder latents with the ground truth features for the SAEs learned with L0=1 and L0=2 compared with the ground-truth SAE in Figure 5. Both these SAEs outperform the ground-truth SAE on variance explained by over 2x despite learning horribly polysemantic latents bearing little resemblance to the underlying true features of the model.
@@ 3.5 Detecting the true L0 using the SAE decoder
Figure 6: Idealized histogram of decoder projections on input activations demonstrating the intuition behind our $n$th decoder projection metric, $s_{n}^{\text{dec}}$. For an arbitrary input, most latents should be non-active and thus have low projection. When SAE latents are monosemantic, we expect non-active latents to have a near-zero projection on arbitrary input activations. However, if SAE latents mix positive and negative components of many underlying features, then those latents will have larger projections on arbitrary inputs that contains those features. By picking an NN less than $h/2$, a smaller $s_{n}^{\text{dec}}$ means latents have smaller projection on arbitrary inputs and thus are more monosemantic.
Figure 1 reveals that the SAE decoder latents contain mixes of underlying features, both when the L0 is too high and also when it is too low. As the SAE approaches the correct L0, each SAE latent has fewer components of multiple true features mixed in, becoming more monosemantic. Thus, we expect that when the SAE is at the correct L0, most latents should have near zero projection on arbitrary training inputs, because they usually do not contain the feature being tracked by that latent. If we are far from the correct L0, then SAE latents contain components of many underlying features, and we expect latents to project more strongly on arbitrary training inputs.
We now define a metric we call $n$th decoder projection score, or $s_{n}^{\text{\text{dec}}}$, that we can use to find the optimal L0 of the SAE. Given SAE inputs $\mathbf{x}\in\mathbb{R}^{b\times d}$ where $b$ is the batch size and $d$ is the input dimension, we first compute the decoder projections for all latents:
$\mathbf{Z}=(\mathbf{x}-\mathbf{b}_{\text{\text{dec}}})\mathbf{W}_{\text{\text{dec}}}^{\top}\in\mathbb{R}^{b\times h}$ (4)
where $\mathbf{b}_{\text{\text{dec}}}\in\mathbb{R}^{d}$ is the decoder bias and $\mathbf{W}_{\text{\text{dec}}}\in\mathbb{R}^{d\times h}$ is the decoder weight matrix with $h$ latent dimensions. To aggregate across the batch, we flatten $\mathbf{Z}$ to obtain $\mathbf{z}\in\mathbb{R}^{bh}$ and sort these values in descending order to get $\mathbf{z}_{\downarrow}$. The $n$th decoder projection is then defined as:
$s_{n}^{\text{\text{dec}}}=\mathbf{z}_{\downarrow}[n\cdot b]$ (5)
@@
Figure 7: $n$th decoder projection vs SAE L0 for $n=12$ (left) and $n=18$ (right) on our toy model SAEs. The true L0, 11, is marked by a dotted line on the plots. Both settings of $n$ are minimized at the true L0, but the slopes of the metric change depending on N. The shaded area is 1 stdev.
We calculate $s_{n}^{\text{\text{dec}}}$ for $n=12$ and $n=18$, varying SAE L0 from 2 to 25 with 5 seeds per L0 in Figure 7. The metric is minimized at the true L0, 11, in both cases, although the shape changes depending on $n$. In both cases, the slope of $s_{n}^{\text{\text{dec}}}$ is flat when L0 is slightly higher than the true L0.
Alternative metric: decoder pairwise cosine similarity
We find that $s_{n}^{\text{\text{dec}}}$ is not the only metric that works to locate the ideal L0. Taking the mean cosine similarity between every latent in the decoder also works well. We focus on $s_{n}^{\text{\text{dec}}}$ as it allows for more interesting analysis, but we do not feel either metric is inherently better than the other, and recommend practitioners experiment with both. We explore this alternative metric further in Appendix A.5.
3.6 JumpReLU SAE experiments
So far, we have only investigated BatchTopK SAES due to their ease of setting L0. We now validate that these same conclusions apply to JumpReLU SAEs. We train JumpReLU saes with a range of $\lambda_{s}$ to control the sparsity of the SAEs. We show plots of $\lambda_{s}$ vs L0 $n$th decoder projection vs L0 for these SAEs in Figure 8. We see that $n$th decoder projection vs L0 broadly follows the same pattern as we saw for BatchTopK SAEs, and is minimized at and slightly above the correct L0.
Figure 8: (left) L0 coefficient $\lambda_{s}$ vs L0 for JumpReLU SAEs. (right) $n$th decoder projection vs L0 for JumpReLU SAEs. The true L0, 11, is marked by a dotted line on the plot.
Interestingly, we see that the L0 does not change linearly with $\lambda_{s}$, but instead “sticks” near the correct L0. This is a testament to Anthropic’s JumpReLU SAE training method (Conerly et al., 2025), as a wide range of sparsity coefficients $\lambda_{s}$ cause the SAE to naturally find the correct L0.